OpenStudy (anonymous):

If the density of a certain spherical atomic nucleus is 1.0 x 10^14 g cm^-3 and its mass is 2.0 x 10^-23 g , what is its radius in cm? Any help?

OpenStudy (anonymous):

@abb0t can you help?

OpenStudy (abb0t):

density has units of \(\sf \frac{mass}{volume}\), which means, you have: \(\sf \color{red}{2.0 \times 10^{-23}~g \times\frac{cm^3}{1.0 \times 10^{14}~ g} }\)